OpenStudy (anonymous):

In a factory that manufactures tires, a machine responsible for molding the tire has a failure rate of 0.2%. If 1,000 tires are produced in a day, of which 6 are faulty, what is the difference between the experimental probability and the theoretical probability? 0.4% 0.2% 0.6% 1%

OpenStudy (dan815):

0.2% of 1000 is?

OpenStudy (dan815):

0.2% = 0.2/100 = 2/1000

OpenStudy (dan815):

0.2% of 1000 means 2/1000 * 1000 = 2

OpenStudy (dan815):

you were expected 2, but you say 6, what percentage is 6 out of 1000?

OpenStudy (dan815):

6/1000=0.6/100= 0.6%

OpenStudy (dan815):

they asked for the difference between these 2, so what is the answer?

OpenStudy (anonymous):

so I need to find out the difference between 0.6% and 2?

OpenStudy (anonymous):

I'm confused on what answers I need to find the difference

OpenStudy (dan815):

the other one is 0.2%

OpenStudy (dan815):

they want the difference between the actual % and the experimental %

OpenStudy (dan815):

do you understand percentages?

OpenStudy (dan815):

percentage means to take a number and rewrite as a fraction with a denominator 100

OpenStudy (anonymous):

yes sorry I was just confused

OpenStudy (dan815):

its fine, ask anything

OpenStudy (anonymous):

so the difference between 0.2 and 0.6 is 0.4?

OpenStudy (anonymous):

A is the correct answer. thanks for the help